# Contrastive Representations for Temporal Reasoning

### v1

@@ 1 Introduction
Machine learning has achieved remarkable progress in vision (Radford et al., , 2021), control (Eysenbach et al., , 2022), and language modeling (Yin et al., , 2024; Jiang et al., , 2024). However, it still falls short on tasks that require structured, combinatorial reasoning. Even relatively simple problems, such as planning in puzzles or verifying symbolic constraints, remain challenging for end-to-end learning systems Valmeekam et al., (2023); Ma et al., (2024). The best methods for solving these problems use computationally expensive search algorithms, such as A* or Best First Search (BestFS) Hart et al., (1968).
@@
This work centers on the question: Can we learn representations that reduce or eliminate the need for search in combinatorial reasoning tasks? To see how good representations can reduce test-time search, consider checking whether a graph has an Euler path. This is true if and only if the graph has exactly zero or two vertices of odd degree. A representation that encodes vertex degrees reduces the task to a simple check—eliminating the need for graph traversal. We approach our question by leveraging temporal contrastive learning (Oord et al., , 2018; Sermanet et al., , 2018; Eysenbach et al., , 2022, 2024; Myers et al., , 2024). These self-supervised techniques are designed to acquire compact, structured representations that capture the problem’s temporal dynamics, enabling efficient planning directly within the latent space.
While (Contrastive Learning)CL has shown promise in control tasks Srinivas et al., (2020); Bortkiewicz et al., (2024), particularly through methods like Contrastive Reinforcement Learning (CRL)Eysenbach et al., (2022), we observe that its effectiveness in combinatorial domains is significantly limited. Specifically, we identify a critical failure mode where contrastive representations overfit to instance-specific context, instead of reflecting environment dynamics. Consequently, models fail to adequately capture the temporal structure that is vital for effective decision-making. This failure mode — such as when the model overfits to wall layouts in Sokoban (Section 4.1) — manifests as a collapse of trajectory representations into small, disconnected clusters, as illustrated in Figure 1 (left).
To solve this, we introduce Contrastive Representations for Temporal Reasoning (CRTR), a simple, theoretically grounded CL method that uses in-trajectory negatives. By design, CRTR forces the model to distinguish between temporally distant states within the same episode. This mechanism prevents it from exploiting irrelevant context — such as visual or layout cues — and instead encourages learning temporally meaningful embeddings that reflect the problem’s relevant dynamics. This echoes recent findings in neuroscience, where hippocampal representations of overlapping routes diverge during learning despite visual similarity Chanales et al., (2017). Our approach similarly prioritizes temporally meaningful structure over reliance on visual cues.
@@ 2 Related Work
Contrastive learning is widely used for discovering rich representations from unlabeled data (Chuang et al., , 2020; Chen et al., , 2020; Jaiswal et al., , 2020; Rethmeier and Augenstein, , 2023) to improve learning downstream tasks (Xiao et al., , 2021). Contrastive learning has enabled effective learning of large-scale models in computer vision (Zhang et al., , 2022; Caron et al., , 2021; Radford et al., , 2021; Liu et al., , 2024) and language (Srivastava et al., , 2023; Gao et al., , 2023). Contrastive learning acquires representations by pulling representations of similar data points, i.e., ones that belong to the same underlying concept, closer together and pushing dissimilar ones further apart in the representation space (Wang and Isola, , 2020). This idea is reflected in various contrastive objectives, including Triplet Loss (Hoffer and Ailon, , 2014), NCE (Gutmann and Hyvärinen, , 2012), InfoNCE (Sohn, , 2016), and conditional InfoNCE (CCE) (Ma et al., , 2022). Our work is most closely related to prior work that uses contrastive learning to obtain representations of time series data (Oord et al., , 2018; Sermanet et al., , 2018; Eysenbach et al., , 2022; Assran et al., , 2023).
@@
Self-supervised contrastive learning has also been applied to sequential (or temporal) problems, including goal-conditioned RL (Eysenbach et al., , 2022; Venkattaramanujam et al., , 2019; Myers et al., , 2024), skill-learning algorithms (Park et al., , 2023; Zheng et al., , 2025; Eysenbach et al., , 2018), and exploration methods (Guo et al., , 2022). Contrastive representations also excel at symbolic reasoning for simple mathematical problems (Poesia et al., , 2021). Most temporal-based contrastive algorithms are based on optimizing the InfoNCE objective (Sohn, , 2016) to distinguish real future states in the trajectory from random states in other trajectories. Interestingly, Eysenbach et al., (2024) demonstrate that intermediate state representations can be effectively approximated via linear interpolation between the initial and final embeddings. Based on these findings, we hypothesize that such representations might also reduce or eliminate the need for search in complex combinatorial problems.
@@
Combinatorial environments are characterized by discrete, compact observations that represent exponentially large configuration spaces, often associated with NP-complete problems Karp, (1972). Recent RL advancements address these challenges using neural networks to learn efficient strategies, including policy-based heuristics Mazyavkina et al., (2021); Bello et al., (2016), graph neural networks for structural exploitation Cappart et al., (2021); Kool et al., (2019), and imitation learning with expert demonstrations Silver et al., (2016). Our work shows that effective search-facilitating representations can be learned from suboptimal data, without relying on expert demonstrations.
@@
Efficient planning in complex environments can be achieved by learning state representations that reflect the underlying structure of the problem. Prior work on world models Ha and Schmidhuber, (2018); Hafner et al., (2023) shows that compact representations of high-dimensional environments are critical for agent performance. Another line of research Schrittwieser et al., (2020); Fang et al., (2022) focuses on learning representations that retain only the features relevant for planning. In robotic settings, similar approaches train latent representations to support movement and decision-making Ichter and Pavone, (2019); Fang et al., (2022). We build on the approach of Eysenbach et al., (2022), which formulates goal-conditioned reinforcement learning as a representation learning problem, with the key distinction that we focus on combinatorial reasoning tasks.
@@ 3 Preliminaries
We employ a contrastive reinforcement learning (CRL) method (Eysenbach et al., , 2022) to train a critic $f(s,g)$, which estimates the similarity between the current state $s$ and future state $g$ (goal). The critic uses a single encoder network $\psi$ to generate both representations of the state $\psi(s)$ and the goal $\psi(g)$. Critic’s output measures a similarity between these representations with a metric $f_{\psi}(s,g)=\|\psi(s)-\psi(g)\|$ that reflects the closeness of the states. For details, see Appendix C. To train the critic, we construct a batch $\mathcal{B}$ by sampling $n$ random trajectories from the dataset. For each trajectory, we select a state $s_{i}$ uniformly and draw a goal $g_{i}$ using a $\operatorname{Geom}(1-\gamma)$ distribution over future states. Negative pairs consist of state-goal pairs $(s_{i},g_{j})$ with goals sampled from different trajectories. We use the critic’s outputs in the InfoNCE objective (Sohn, , 2016), following prior work in CRL (Eysenbach et al., , 2022, 2021; Zheng et al., , 2023, 2024; Myers et al., , 2024; Bortkiewicz et al., , 2024). This objective provides a lower bound on the mutual information between state and goal representations:
@@
Several prior works use explicit search algorithms solving complex environments Silver et al., (2016); Brown et al., (2020); Yonetani et al., (2021); Orseau et al., (2018). In our study, we focus on the Best-First Search (BestFS) Pearl, (1984). BestFS builds the search tree by greedily expanding nodes with the highest heuristic estimates, hence targeting paths that are most likely to lead to the goal. In our approach, BestFS uses the representations of the current state and goal, and relies on the critic to evaluate the distances between neighboring states and the goal. While not ensuring optimality, BestFS provides a simple yet effective strategy for navigating complex search spaces. The pseudocode for BestFS is outlined in Appendix B. In our work, we use distances in the latent space as the heuristic, as detailed in Section 3.
@@ 4.1 Failure of Naïve Temporal Contrastive Learning in Combinatorial Domains
A straightforward approach to learning representations $\phi(s)$ is to use temporal contrastive learning (Oord et al., , 2018; Sermanet et al., , 2018) (outlined in Sec. 3): positive pairs are sampled from nearby states within the same trajectory, and negatives from different trajectories. However, applying this approach to a common benchmark (Sokoban) reveals an important failure mode. This section introduces this failure mode and its mathematical underpinnings, Sec. 4.2 introduces an idealized algorithm for fixing this failure mode, and Sec. 4.3 turns this idealized algorithm into a practical one that we use for our experiments in Sec. 5.
@@ 4.2 Learning Representations that Ignore Context: An Idealized Algorithm
Mathematically, this alternative objective is a lower bound on the conditional mutual information $I(X;X_{+}\mid C)$ (Ma et al., , 2022), which can be decomposed using the chain rule for mutual information:
@@
The term $I(X;X_{+})$ is what is usually maximized by temporal contrastive learning. The second term, $I(X_{+};C)$, is akin to adversarial feature learning (Li et al., , 2018; Donahue et al., , 2016; Xie et al., , 2017), which aims to learn representations that do not retain certain pieces of information. Our method achieves a similar effect without the need for adversarial optimization. This context-invariance objective parallels mechanisms observed in biological systems, where overlapping episodic memories are actively decorrelated to reduce interference Chanales et al., (2017).
@@ 5.1 Experimental Setup
We evaluate all methods on five challenging combinatorial reasoning tasks: Sokoban (Dor and Zwick, , 1999), Rubik’s Cube, N-Puzzle (Johnson and Story, , 1879), Lights Out (Anderson and Feil, , 1998), and Digit Jumper (Bagatella et al., , 2021). Most of these are NP-hard Demaine et al., (2018); Culberson, (1998); Ratner and Warmuth, (1986) and serve as standard RL benchmarks Agostinelli et al., (2019); Racanière et al., (2017); Zawalski et al., (2024). Sokoban is a grid-based puzzle where an agent pushes boxes to targets while avoiding irreversible states. Rubik’s Cube requires aligning each face of a 3D cube to a single color. N-Puzzle involves sliding tiles within a $4\times 4$ grid to reach a goal configuration. Lights Out is a toggle-based puzzle aiming to switch all cells to an off state. Digit Jumper is a grid game where each cell indicates the jump length from that position. See Appendix A for full environment details.
@@
We compare against three baselines. The contrastive baseline performs temporal contrastive learning Sermanet et al., (2018); Oord et al., (2018); Eysenbach et al., (2022), training representations without in-trajectory negatives. We will refer to this baseline as contrastive RL (CRL) (Eysenbach et al., , 2022). The supervised baseline (Czechowski et al., , 2021; Zawalski et al., , 2023) predicts state distances using a value network trained via imitation on demonstrations. Finally, the DeepCubeA Agostinelli et al., (2019) baseline learns a value function via iterative one-step lookahead. As a lower bound, we also evaluate the performance of representations from a randomly initialized network. For fairness, the CRL and supervised baselines use the same architecture as CRTR.
@@ 5.2 Context-Free Representations for Combinatorial Reasoning
Figure 2: CRTR performs well in all the evaluated domains. Success rate as a function of search budget across five domains. CRTR compared to baselines: CRL Eysenbach et al., (2022), Supervised Czechowski et al., (2021) and DeepCubeA Agostinelli et al., (2019). Results are averaged over 5 seeds; shaded regions indicate standard error. For the Rubik’s Cube, both the supervised and random baselines achieve a success rate of zero for all search budgets.
@@
Our third experiment compares CRTR to supervised approaches Agostinelli et al., (2019); Czechowski et al., (2021) for solving combinatorial problems. Again, see Fig. 2. CRTR ranks among the top-performing methods in each environment and is strictly the best in two environments. In contrast, the supervised baselines perform much worse in Rubik’s Cube and Lights Out. We conjecture that the improved performance of CRTR come from how it represents values as distance between learned representations, rather than a number output by a monolithic neural network.
@@ 6 Conclusion
Another way of reading this paper is as a story about the value of metric embeddings. Image classifiers automatically identify patterns and structures in images (arrays of pixels), mapping images to representations so that semantically-similar images are mapped to similar representations. Our representation learning method does the same for combinatorial reasoning problems, automatically identifying patterns and structures of observations so that the simplest possible decision rule (greedy action selection) can solve some of the most complex reasoning problems (e.g., Sokoban in P-SPACE complete (Culberson, , 1998)). The success of these representations raises the question of how much reasoning can be done in the representation space. Can we think about compositional reasoning as vector addition? Can we think of length generalization as rays in representation space? And, perhaps most alluringly, can we exploit representational geometry to use the solutions to simple problems to find solutions to hard problems, by using local rules to provide global structure on the space of representations?
@@ Appendix A Environments
Sokoban is a well-known puzzle game in which a player pushes boxes onto designated goal positions within a confined grid. It is known to be hard from a computational complexity perspective. Solving it requires reasoning over a vast number of possible move sequences, making it a standard benchmark for both classical planning algorithms and modern deep learning approaches Dor and Zwick, (1999). Solving Sokoban requires balancing efficient search with long-term planning. In our experiments, we use 12×12 boards with four boxes.
@@
The Rubik’s Cube is a 3D combinatorial puzzle with over $4.3\times 10^{19}$ possible configurations, making it an ideal testbed for algorithms tackling massive search spaces. Solving the Rubik’s Cube requires sophisticated reasoning and planning, as well as the ability to efficiently navigate high-dimensional state spaces. Recent advances in using neural networks for solving this puzzle, such as Agostinelli et al., (2019), highlight the potential of deep learning in handling such computationally challenging tasks.
@@ Appendix B Best-First Search
Algorithm 2 Best-First Search (Hart et al., , 1968)
@@ Appendix C Training Details
For Sokoban, we use trajectories provided by Czechowski et al., (2021) and train on a dataset of $10^{5}$ trajectories. For 15-Puzzle, Rubik’s Cube, and Lights Out, we generate training trajectories by applying a policy that performs $n$ random actions, where $n$ is set to 150, 21, and 49, respectively. In the case of 15-Puzzle, we additionally remove single-step cycles from the dataset to improve data efficiency. For Digit Jumper, we generate training data by sampling a random path from the upper-left corner to the bottom-right corner on a standard $20\times 20$ grid. All grid cells not required for this path are filled by sampling uniformly from the set $1,\dots,6$}. The network for Digit Jumper typically converges after a few hours of training, so we train until convergence is observed. For Sokoban, Rubik’s Cube, Lights Out, and 15-Puzzle, we adopt an unlimited data setup and train all models for two days. This results in the models performing approximately $8\times 10^{6}$ gradient updates for Rubik’s Cube, $7\times 10^{6}$ for 15-Puzzle, and $9\times 10^{6}$ for Lights Out.
@@
We adopt the network architecture proposed by Nauman et al., (2025), using 8 layers with a hidden size of 512 and a representation dimension of 64. This configuration was found to yield optimal performance for the contrastive baseline on Rubik’s Cube, as illustrated in Figure 7. We observed that this architecture performs well in all environments except for two cases:

### v2

@@ 1 Introduction
Machine learning has achieved remarkable progress in vision [51], control [21], and language modeling [65, 33]. However, it still falls short on tasks that require structured, combinatorial reasoning. Even relatively simple problems, such as planning in puzzles or verifying symbolic constraints, remain challenging for end-to-end learning systems [60, 40]. The best methods for solving these problems use computationally expensive search algorithms, such as A* or Best First Search (BestFS) [29].
@@
This work centers on the question: Can we learn representations that reduce or eliminate the need for search in combinatorial reasoning tasks? To see how good representations can reduce test-time search, consider checking whether a graph has an Euler path. This is true if and only if the graph has exactly zero or two vertices of odd degree. A representation that encodes vertex degrees reduces the task to a simple check—eliminating the need for graph traversal. We approach our question by leveraging temporal contrastive learning [45, 55, 21, 19, 43]. These self-supervised techniques are designed to acquire compact, structured representations that capture the problem’s temporal dynamics, enabling efficient planning directly within the latent space.
While (Contrastive Learning)CL has shown promise in control tasks [58, 6], particularly through methods like Contrastive Reinforcement Learning (CRL)[21], we observe that its effectiveness in combinatorial domains is significantly limited. Specifically, we identify a critical failure mode where contrastive representations overfit to instance-specific context, instead of reflecting environment dynamics. Consequently, models fail to adequately capture the temporal structure that is vital for effective decision-making. This failure mode — such as when the model overfits to wall layouts in Sokoban (Section 4.1) — manifests as a collapse of trajectory representations into small, disconnected clusters, as illustrated in Figure 1 (left).
To solve this, we introduce Contrastive Representations for Temporal Reasoning (CRTR), a simple, theoretically grounded CL method that uses in-trajectory negatives. By design, CRTR forces the model to distinguish between temporally distant states within the same episode. This mechanism prevents it from exploiting irrelevant context — such as visual or layout cues — and instead encourages learning temporally meaningful embeddings that reflect the problem’s relevant dynamics. This echoes recent findings in neuroscience, where hippocampal representations of overlapping routes diverge during learning despite visual similarity [10]. Our approach similarly prioritizes temporally meaningful structure over reliance on visual cues.
@@ 2 Related Work
Contrastive learning is widely used for discovering rich representations from unlabeled data [12, 11, 32, 53] to improve learning downstream tasks [63]. Contrastive learning has enabled effective learning of large-scale models in computer vision [69, 9, 51, 39] and language [59, 24]. Contrastive learning acquires representations by pulling representations of similar data points, i.e., ones that belong to the same underlying concept, closer together and pushing dissimilar ones further apart in the representation space [62]. This idea is reflected in various contrastive objectives, including Triplet Loss [30], NCE [26], InfoNCE [57], and conditional InfoNCE (CCE) [41]. Our work is most closely related to prior work that uses contrastive learning to obtain representations of time series data [45, 55, 21, 3].
@@
Self-supervised contrastive learning has also been applied to sequential (or temporal) problems, including goal-conditioned RL [21, 61, 43], skill-learning algorithms [47, 72, 18], and exploration methods [25]. Contrastive representations also excel at symbolic reasoning for simple mathematical problems [49]. Most temporal-based contrastive algorithms are based on optimizing the InfoNCE objective [57] to distinguish real future states in the trajectory from random states in other trajectories. Interestingly, Eysenbach et al., [19] demonstrate that intermediate state representations can be effectively approximated via linear interpolation between the initial and final embeddings. Based on these findings, we hypothesize that such representations might also reduce or eliminate the need for search in complex combinatorial problems.
@@
Combinatorial environments are characterized by discrete, compact observations that represent exponentially large configuration spaces, often associated with NP-complete problems [35]. Recent RL advancements address these challenges using neural networks to learn efficient strategies, including policy-based heuristics [42, 5], graph neural networks for structural exploitation [8, 36], and imitation learning with expert demonstrations [56]. Our work shows that effective search-facilitating representations can be learned from suboptimal data, without relying on expert demonstrations.
@@
Efficient planning in complex environments can be achieved by learning state representations that reflect the underlying structure of the problem. Prior work on world models [27, 28] shows that compact representations of high-dimensional environments are critical for agent performance. Another line of research [54, 22] focuses on learning representations that retain only the features relevant for planning. In robotic settings, similar approaches train latent representations to support movement and decision-making [31, 22]. We build on the approach of Eysenbach et al., [21], which formulates goal-conditioned reinforcement learning as a representation learning problem, with the key distinction that we focus on combinatorial reasoning tasks.
@@ 3 Preliminaries
We employ a contrastive reinforcement learning (CRL) method [21] to train a critic $f(s,g)$, which estimates the similarity between the current state $s$ and future state $g$ (goal). The critic uses a single encoder network $\psi$ to generate both representations of the state $\psi(s)$ and the goal $\psi(g)$. Critic’s output measures a similarity between these representations with a metric $f_{\psi}(s,g)=\|\psi(s)-\psi(g)\|$ that reflects the closeness of the states. For details, see Appendix C. To train the critic, we construct a batch $\mathcal{B}$ by sampling $n$ random trajectories from the dataset. For each trajectory, we select a state $s_{i}$ uniformly and draw a goal $g_{i}$ using a $\operatorname{Geom}(1-\gamma)$ distribution over future states. Negative pairs consist of state-goal pairs $(s_{i},g_{j})$ with goals sampled from different trajectories. We use the critic’s outputs in the InfoNCE objective [57], following prior work in CRL [21, 20, 71, 70, 43, 6]. This objective provides a lower bound on the mutual information between state and goal representations:
@@
Several prior works use explicit search algorithms solving complex environments [56, 7, 66, 46]. In our study, we focus on the Best-First Search (BestFS) [48]. BestFS builds the search tree by greedily expanding nodes with the highest heuristic estimates, hence targeting paths that are most likely to lead to the goal. In our approach, BestFS uses the representations of the current state and goal, and relies on the critic to evaluate the distances between neighboring states and the goal. While not ensuring optimality, BestFS provides a simple yet effective strategy for navigating complex search spaces. The pseudocode for BestFS is outlined in Appendix B. In our work, we use distances in the latent space as the heuristic, as detailed in Section 3.
@@ 4.1 Failure of Naïve Temporal Contrastive Learning in Combinatorial Domains
A straightforward approach to learning representations $\phi(s)$ is to use temporal contrastive learning [45, 55] (outlined in Sec. 3): positive pairs are sampled from nearby states within the same trajectory, and negatives from different trajectories. However, applying this approach to a common benchmark (Sokoban) reveals an important failure mode. This section introduces this failure mode and its mathematical underpinnings, Sec. 4.2 introduces an idealized algorithm for fixing this failure mode, and Sec. 4.3 turns this idealized algorithm into a practical one that we use for our experiments in Sec. 5.
@@ 4.2 Learning Representations that Ignore Context: An Idealized Algorithm
Mathematically, this alternative objective is a lower bound on the conditional mutual information $I(X;X_{+}\mid C)$ [41], which can be decomposed using the chain rule for mutual information:
@@
The term $I(X;X_{+})$ is what is usually maximized by temporal contrastive learning. The second term, $I(X_{+};C)$, is akin to adversarial feature learning [38, 16, 64], which aims to learn representations that do not retain certain pieces of information. Our method achieves a similar effect without the need for adversarial optimization. This context-invariance objective parallels mechanisms observed in biological systems, where overlapping episodic memories are actively decorrelated to reduce interference [10].
@@ 5.1 Experimental Setup
We evaluate all methods on five challenging combinatorial reasoning tasks: Sokoban [17], Rubik’s Cube, N-Puzzle [34], Lights Out [2], and Digit Jumper [4]. Most of these are NP-hard [15, 13, 52] and serve as standard RL benchmarks [1, 50, 68]. Sokoban is a grid-based puzzle where an agent pushes boxes to targets while avoiding irreversible states. Rubik’s Cube requires aligning each face of a 3D cube to a single color. N-Puzzle involves sliding tiles within a $4\times 4$ grid to reach a goal configuration. Lights Out is a toggle-based puzzle aiming to switch all cells to an off state. Digit Jumper is a grid game where each cell indicates the jump length from that position. See Appendix A for full environment details.
@@
We compare against three baselines. The contrastive baseline performs temporal contrastive learning [55, 45, 21], training representations without in-trajectory negatives. We will refer to this baseline as contrastive RL (CRL) [21]. The supervised baseline [14, 67] predicts state distances using a value network trained via imitation on demonstrations. Finally, the DeepCubeA [1] baseline learns a value function via iterative one-step lookahead. As a lower bound, we also evaluate the performance of representations from a randomly initialized network. For fairness, the CRL and supervised baselines use the same architecture as CRTR.
@@ 5.2 Context-Free Representations for Combinatorial Reasoning
Figure 2: CRTR performs well in all the evaluated domains. Success rate as a function of search budget across five domains. CRTR compared to baselines: CRL [21], Supervised [14] and DeepCubeA [1]. Results are averaged over 5 seeds; shaded regions indicate standard error. For the Rubik’s Cube, both the supervised and random baselines achieve a success rate of zero for all search budgets.
@@
Our third experiment compares CRTR to supervised approaches [1, 14] for solving combinatorial problems. Again, see Fig. 2. CRTR ranks among the top-performing methods in each environment and is strictly the best in two environments. In contrast, the supervised baselines perform much worse in Rubik’s Cube and Lights Out. We conjecture that the improved performance of CRTR come from how it represents values as distance between learned representations, rather than a number output by a monolithic neural network.
@@ 6 Conclusion
Another way of reading this paper is as a story about the value of metric embeddings. Image classifiers automatically identify patterns and structures in images (arrays of pixels), mapping images to representations so that semantically-similar images are mapped to similar representations. Our representation learning method does the same for combinatorial reasoning problems, automatically identifying patterns and structures of observations so that the simplest possible decision rule (greedy action selection) can solve some of the most complex reasoning problems (e.g., Sokoban in P-SPACE complete [13]). The success of these representations raises the question of how much reasoning can be done in the representation space. Can we think about compositional reasoning as vector addition? Can we think of length generalization as rays in representation space? And, perhaps most alluringly, can we exploit representational geometry to use the solutions to simple problems to find solutions to hard problems, by using local rules to provide global structure on the space of representations?
@@ Appendix A Environments
Sokoban is a well-known puzzle game in which a player pushes boxes onto designated goal positions within a confined grid. It is known to be hard from a computational complexity perspective. Solving it requires reasoning over a vast number of possible move sequences, making it a standard benchmark for both classical planning algorithms and modern deep learning approaches [17]. Solving Sokoban requires balancing efficient search with long-term planning. In our experiments, we use 12×12 boards with four boxes.
@@
The Rubik’s Cube is a 3D combinatorial puzzle with over $4.3\times 10^{19}$ possible configurations, making it an ideal testbed for algorithms tackling massive search spaces. Solving the Rubik’s Cube requires sophisticated reasoning and planning, as well as the ability to efficiently navigate high-dimensional state spaces. Recent advances in using neural networks for solving this puzzle, such as [1], highlight the potential of deep learning in handling such computationally challenging tasks.
@@ Appendix B Best-First Search
Algorithm 2 Best-First Search [29]
@@ Appendix C Training Details
For Sokoban, we use trajectories provided by Czechowski et al., [14] and train on a dataset of $10^{5}$ trajectories. For 15-Puzzle, Rubik’s Cube, and Lights Out, we generate training trajectories by applying a policy that performs $n$ random actions, where $n$ is set to 150, 21, and 49, respectively. In the case of 15-Puzzle, we additionally remove single-step cycles from the dataset to improve data efficiency. For Digit Jumper, we generate training data by sampling a random path from the upper-left corner to the bottom-right corner on a standard $20\times 20$ grid. All grid cells not required for this path are filled by sampling uniformly from the set $1,\dots,6$}. The network for Digit Jumper typically converges after a few hours of training, so we train until convergence is observed. For Sokoban, Rubik’s Cube, Lights Out, and 15-Puzzle, we adopt an unlimited data setup and train all models for two days. This results in the models performing approximately $8\times 10^{6}$ gradient updates for Rubik’s Cube, $7\times 10^{6}$ for 15-Puzzle, and $9\times 10^{6}$ for Lights Out.
@@
We adopt the network architecture proposed by Nauman et al., [44], using 8 layers with a hidden size of 512 and a representation dimension of 64. This configuration was found to yield optimal performance for the contrastive baseline on Rubik’s Cube, as illustrated in Figure 7. We observed that this architecture performs well in all environments except for two cases: